find the (GCF) greatest common factor of each set of numbers:
1. 12 and 22
2. 7 and 21
3. 24 and 48

Answer by rapaljer(4551)

Factor each number into primes:
12=3*4
12=2^2*3

22= 2*11

GCF = Product of all common factors = 2

7= 7 (Prime)
21 = 3*7

GCF = Product of all common factors = 7

24 and 48
Because 24 divides into 48 evenly, this means that 24 is the GCF without going to all the trouble of factoring the numbers into primes. If you did factor, the result would be 24.

R^2 at SCC